# A Comprehensive Study on GDPR-Oriented Analysis of Privacy Policies: Taxonomy, Corpus and GDPR Concept Classifiers

### v1

@@ Abstract
Machine learning based classifiers that take a privacy policy as the input and predict relevant concepts are useful in different applications such as (semi-)automated compliance analysis against requirements of the EU GDPR. In all past studies, such classifiers produce a concept label per segment (e.g., sentence or paragraph) and their performances were evaluated by using a dataset of labeled segments without considering the privacy policy they belong to. However, such an approach could overestimate the performance in real-world settings, where all segments in a new privacy policy are supposed to be unseen. Additionally, we also observed other research gaps, including the lack of a more complete GDPR taxonomy and the less consideration of hierarchical information in privacy policies. To fill such research gaps, we developed a more complete GDPR taxonomy, created the first corpus of labeled privacy policies with hierarchical information, and conducted the most comprehensive performance evaluation of GDPR concept classifiers for privacy policies. Our work leads to multiple novel findings, including the confirmed inappropriateness of splitting training and test sets at the segment level, the benefits of considering hierarchical information, and the limitations of the “one size fits all” approach, and the significance of testing cross-corpus generalizability.
Index Terms:
GDPR taxonomy, privacy policy corpus, legal compliance, concept classifier
@@ I Introduction
The importance of privacy policies for both data subjects and data controllers means that GDPR-oriented analyses of such legal documents can be very useful. It can help enhance data subjects’ awareness on important data protection issues, and also help data controllers refine their privacy policy to be more legally compliant to the GDPR. Such an analysis can obviously be done qualitatively by experienced GDPR experts, but more automated analysis is preferred in many applications to help reduce the time and costs spent. As reviewed in Section II of this paper, many researchers have investigated the automation of analysis of privacy policies, but they all evaluated their concept classifiers by splitting the training and testing sets at the segment level without considering which segment belongs to which privacy policy. Such an approach means that segments used in both the training and testing sets can cover all privacy policies, while in real-world applications the classifiers should be tested on segments from unseen privacy policies. Therefore, a classifier tested using the segment-level treatment could perform worse in real-world settings since its training and testing sets may have not covered all segments of an unseen privacy policy. In addition, we also observed other research gaps in the literature, including the lack of a more complete GDPR taxonomy and the less consideration of hierarchical information in privacy policies for developing GDPR-oriented concept classifiers.
In this paper, we fill all the above research gaps by presenting the most comprehensive comparative study on the performances of GDPR concept classifiers. Our study involved the development of a more complete GDPR taxonomy, the first privacy policy corpus covering hierarchical information, and a comprehensive performance analysis of many different GDPR concept classifiers based on a new document-level training-testing set splitting method, two architectures of hierarchical classifiers, and many different sets of features including some covering contextual information. More details about our contributions are summarized as follows (more significant contributions appearing first).
1) We propose a document-level performance evaluation framework for GDPR concept classifiers, and obtained the actual performance of the classifier in real-world scenarios. Different from previous segment-level evaluation, we divide the training and testing sets by documents so that no segments in the testing set are included in the training process. By comparing the performance results at the document and segment levels, we showed that the performances of GDPR concept classifiers reported in past studies are indeed significantly overestimated, so cannot reflect their real-world performances. This indicates the need for all future studies to follow the document-level performance framework.
2) We conduct a comparative study on GDPR concept classifiers (the most comprehensive one to the best of our knowledge). Our study goes beyond past research in multiple ways: i) considering contextual features based on the hierarchical nature of a privacy policy (derived from the parent node and the immediate sibling nodes); ii) comparing two architectures of hierarchical classifiers – local classifier per node (LCN) and local classifier per parent node (LCPN), not just LCPN in past work; iii) obtaining a range of new insights based on extensive experiments, such as evidence against the “one type of classifier fits all” idea followed in past work and the error propagation issue of LCPN classifiers.
3) We constructed a new and the first fully hierarchically encoded GDPR-oriented privacy policy corpus GoPPC-150 and the new framework for extending this new corpus and constructing other new corpora. GoPPC-150 includes 150 privacy policies collected from Alexa.com top websites. It includes expert-annotated GDPR concept labels in a hierarchical manner following a newly extended GDPR privacy policy taxonomy (see below), and is structured at the document-level. The inclusion of explicitly encoded hierarchical information in GoPPC-150 allows more context-aware development of GDPR concept classifiers and other relevant tools. The framework has a high level of automation to minimize human intervention.
@@
4) We propose an extended taxonomy for facilitating GDPR-oriented analysis of privacy policies. This taxonomy is based on a smaller GDPR taxonomy proposed by Torre et al. [7], our own expertise, two important documents from the ICO and the IAPP (International Association of Privacy Professionals), and the work of the W3C Data Privacy Vocabularies and Controls CG (DPVCG). The end result is the most comprehensive GDPR privacy policy taxonomy reported in the literature so far.
@@
To help other researchers reproduce our work reported in this paper, all our source code, data used and newly produced, and all results of our experiments have been uploaded to a GitHub repository available at https://github.com/tp-sh/GDPR_privacy_policies. We will refer to this GitHub repository frequently in the rest of the paper, with a direct URL pointing to the corresponding part of the repository when necessary.
@@
The rest of this paper is structured as follows. Section II discusses related work. Section III shows the initial experimental evidence regarding the significantly reduced performances of selected concept classifiers previously trained at the segment level, when they are applied to unseen privacy policies. Section IV explains the more complete GDPR taxonomy and the new hierarchical corpus we developed, including the methods we used for their development. Section V introduces the comprehensive study on the GDPR concept classifiers and the evaluation results using the document-level evaluation framework. Section VI discusses the further research directions of our work and the field, and the last section concludes the paper.
@@ II Related Work
In this section, we introduce related work in five closely related areas: (1) hierarchical multi-label text classification, (2) privacy policy corpora, (3) privacy policy concept classifiers, and (4) analyses of privacy policies.
Hierarchical multi-label text classification (HMTC). Privacy policy is a hierarchical text with multi-label, so the compliance checking of privacy policy can be regarded as an HMTC task. There are many studies on HMTC, in 2011, Silla et al. [8] summarized the three methods of HMTC, which are flat, local and global approaches. To achieve better results, the latter two are more concerned. Koller et al. [9] introduced the first type of local classifier which is proposed to explore the hierarchy by using local information and build multiple local classifiers around it. Following which, a series of local approaches including LCN (a local classifier per node), LCPN (a local classifier per parent node) as well as LCL (a local classifier per level) based methods are proposed [10, 11, 12]. However, local approaches on their own possess high risk of error-propagation. On the other hand, global approaches are designed to build a single classifier to explore hierarchical information globally, thus reducing the overall model size. Most global approaches are generally modified from the flat classification algorithms such as hAnt-Miner [13], Vens et al. [14]) (decision tree based) and GMNBwU [15](Naive Bayes classifier based). Recently, more neural network based HMTC algorithms which combine both local and global approaches are proposed [16, 17]. Huang [18] proposed HARNN, extracting hierarchical information at different label levels and generating document embeddings for local and global classifications. Zhang [19] introduced LA-HCN which employs label-based attention based on the individual label, and the respective embeddings are obtained for global and local classification for better error propagation control. However, when it comes to the compliance checking of privacy policy, few researches have applied global approaches for classification. Most of the classifiers are local [20, 21, 7, 22, 23]. Their specific approaches are shown in following Table I. The tuples in column Taxonomy explain the numbers of levels and nodes of the taxonomies as (3, 96) represents there are 96 nodes of 3-level in the taxonomy. The tuples in column Corpus/Corpora explain the number of samples, annotation levels and be in hierarchical structure or not. For example, (150, 3, Yes) represents that there are 150 privacy policies with annotations of 3-level in the corpus GoPPC-150 and has a hierarchical structure.
TABLE I: Our work compared with related work (* = new public corpus;  = GDPR-specific)
@@
Related Work Taxonomy Features Corpus/Corpora Classifier(s) Analysis
OPP-115 [24] (2, 46) Paragraph2Vec OPP-115* (115, 2, No) Multiple; multi-class; LCPN OPP-115
Polisis [20] No fastText OPP-115 CNN; multi-class; LCPN No
Mustapha et al.’s [21] No XLNet OPP-115 XLNet; multi-class; LCPN No
PrivBERT [23] No RoBERTa OPP-115 CNN; multi-class; LCPN No
Torre et al.’s [25] (3, 69) GloVe Private SVM; binary; LCPN Private corpus
Rahat et al.’s [22] (1, 18) fastText Private CNN; multi-class; LCPN Private corpus
Our Work (3, 96) Multiple Context GoPPC-150* (150, 3, Yes) OPP-115 Multiple; binary + multi-class; LCPN + LCN GoPPC-150 Template-based
@@
Privacy policy corpora. In 2016, Wilson et al. [24] created a public privacy policy corpus called “Online Privacy Policy” (OPP-115) by hiring three legal experts as annotators and it has been used by many other researchers. However, Sarne et al. [26] used unsupervised ML to model topics in privacy policies and found a mismatch between the topics in privacy policies they analyzed and topics covered in OPP-115. Sathyendra et al. [27] constructed a more fine-grained corpus based on OPP-115 with semi-automated annotation, focusing on “opt-out” in privacy policies. Leblanc and Liu [28] built a corpus of fuzzy words and sentences in privacy policies. Zimmeck et al. [29] created the App-350 Privacy Policy Corpus with the goal of checking compliance of behaviors and privacy policies of mobile apps. They selected 350 privacy policies of the most popular apps on Google Play and hired legal experts as annotators. Robaldo et al. [30] used I/O logic formula to code and model various legal documents, expressed legal statements in a logical language, explained various clauses, and demonstrated their work with the GDPR. Muller et al. [31] introduced a privacy policy dataset containing over 18,300 sentences, tagged according to five core privacy policy requirements of the GDPR. Srinath et al. [23] created PrivaSeer, an automatically constructed corpus containing privacy policies of over 1 million English websites, and studied the composition of the corpus, showing readability tests, document similarity, keyword extraction results, and exploring the corpus through topic modeling. Kuznetsov et al. [32] used a technique for identifying URLs of privacy policies of IoT devices to construct a new corpus with 592 privacy policies in 2022.
@@
Privacy policy concept classifiers. Harkous et al. [20] proposed Polisis, a comprehensive framework for enabling multi-dimensional privacy policy analysis. Polisis provides an online service for privacy policies analysis, using a combination of NLP and deep learning (DL) techniques to extract fragments from privacy policies, each containing a set of labels describing data processing behaviors. Tesfay et al. [33] proposed an ML-based approach to classify multiple categories of privacy policy content using pre-defined keywords. Lippi et al. [34] provided 33 metadata types for privacy policies in terms of the GDPR compliance, and provided automatic support for ambiguity detection of privacy policies based on manual rules and machine learning (ML) based methods. A more comprehensive piece of work was done by Torre et al. [7], who proposed 55 metadata types, covering all types identified by Lippi et al. [34], and studied automated detection of privacy policy integrity using an advanced combination of natural language processing (NLP) and ML based on the 55 metadata types. Their more recent work [25] further improved their privacy policy completeness checking framework with the same GDPR taxonomy and classifiers. Srinath et al. [23] constructed a new privacy policy corpus called PrivSeer, built a pre-trained language model called PrivBERT based on PrivSeer, and then developed a number of concept classifiers for privacy policy analysis based on PriVERT. As a whole, the best-performing classifiers are PrivBERT-based ones and those reported in [25], achieving an F1 score above 0.8 for most concept classifiers, and even above 0.9 for some.
@@
Analyses of privacy policies. There are many past studies focusing on the analyses of privacy policies, some of which are GDPR-oriented. Many of such studies are based on machine learning based concept classifiers, as part of a larger automated system. When it comes to the GDPR-oriented analyses, Tom et al. [35] presented a preliminary GDPR model that aims to provide a simple and visual overview for human operators to achieve a better understanding of the relationships between different concepts in the GDPR. It also describes a method of using their proposed model as a tool to develop privacy policies and illustrates how to extract compliance rules. Palmirani and Governatori [36] proposed a proof of concept applicable to the GDPR domain with the aim of detecting or preventing violations of privacy enforcement norms. Bhatia et al. [37] identified incompleteness of privacy policies by representing data practice descriptions as semantic frames. The approach was a grounded analysis to discover which semantic roles corresponding to a data action are needed to construct complete data practice descriptions. Mousavi et al. [38] proposed a tool called KnIGHT, whose innovation lies in the use of semantic similarity between words to associate sentences in a privacy policy with relevant paragraphs in the legal text of the GDPR. Torre et al. [39] proposed a model-based solution for GDPR-compliance analysis, using Unified Modeling Language (UML) and Object Constraint Language (OCL) to build a UML representation of the GDPR. Hamdani et al. [40] conceptualized a framework to implement document-central compliance checking methods in the data supply chain and developed concrete methods to automatically check GDPR-compliance of privacy policies. The other work focus on the commonalities of privacy policies. In order to increase transparency of privacy policies, Zimmeck et al. [41] proposed Privee, a software architecture for analyzing essential policy terms based on crowd sourcing and automatic classification techniques. They implemented Privee as a proof-of-concept web browser extension that retrieves policy analysis results from an online privacy policy repository or, if no such results are available, performs automatic classifications. Targeting privacy policies of websites and mobile apps, Caramujo et al. [42] proposed a domain-specific language and a model transformation approach for specifying privacy policy models. Pullonen et al. [43] proposed a multi-level model as an extension of the business process model and representation to support visualization, analysis, and communication of the privacy policy characteristics of business processes. Ayala-Rivera and Pasquale [44] proposed a model-based approach to help online services understand their data protection obligations under the GDPR. Zimmeck et al. [45] designed and implemented PrivacyFlash Pro, an automated privacy policy generator for iOS apps that leverages static analysis, which identifies code signatures composed of Plist permission strings, framework imports, class instantiations, authorization methods and other evidence that are mapped to privacy practices expressed in privacy policies. Wang et al. [46] proposed PRIVGUARD, a novel system design that reduces human participation required and improves the productivity of the compliance process, which is mainly comprised of two components: (1) PRIVANALYZER, a static analyzer based on abstract interpretation for partly enforcing privacy regulations, and (2) a set of components providing strong security protection on the data throughout its life cycle. Cui et al. [47] proposed POLIGRAPH, automated analysis of the information disclosed in a policy into a knowledge graph based on NLP technology. Although they also combine the context of the current text for classification, they only focus on the data collection part of the privacy policy, and do not consider the structural information of the privacy policy.
III Potential Performance Overestimation of Existing Concept Classifiers
As explained in the Introduction section, all past studies on concept classifiers of privacy policies all followed the segment-level performance evaluation approach. In order to show that such an approach could lead to an overestimation of the performance, we conducted a small experiment to test the performances of some PrivBERT-based concept classifiers with a good performance as reported in [23] [^2] For this experiment, we did not select Torre et al.’s classifiers [25] because of the following reasons: 1) their classifiers are not open sourced; 2) they used various artificial rules so it is harder to generalize their classifiers; and 3) we reproduced their classifiers based on their descriptions in [25] and tested their classifiers’ performances against the new corpus we developed, but all the classifiers performed turned out to be very poor (see the results at https://github.com/tp-sh/GDPR_privacy_policies/blob/main/data/Torre_classifiers_results.csv), indicating that their classifiers may have overfitting problems., by applying them to three privacy policies the first author of the paper manually labeled. These classifiers, which were previously trained and tested at the segment level using the OPP-115 corpus, now needed to make predictions on unseen privacy policies, which are not included in the OPP-115 corpus. The results showed that all classifiers performed much worse than what was reported in [23]. Table II shows the performance comparison of 4 concept classifiers with more than 10 positive samples, and the full results of all the 10 concept classifiers can be found at https://github.com/tp-sh/GDPR_privacy_policies/blob/main/data/opp_results.csv.
TABLE II: Performance of classifier under different evaluation method
OPP-115 Concept Old Evaluation (segment-level) Our Evaluation (document-level)
P R F1 P R F1
FIRST PARTY COLLECTION AND USE 0.913 0.945 0.929 51/70 (-18.4%) 51/58 (-6.5%) 0.797 (-13.2%)
THIRD PARTY SHARING AND COLLECTION 0.915 0.945 0.930 24/58 (-49.7%) 24/25 (+1.5%) 0.578 (-35.2%)
USER CHOICE/CONTROL 0.906 0.828 0.865 11/21 (-38.2%) 11/27 (-42.1%) 0.458 (-40.7%)
INTERNATIONAL AND SPECIFIC AUDIENCES 0.960 0.947 0.954 25/28 (-6.7%) 25/35 (-23.3%) 0.794 (-16.0%)
The results indicate that the PrivBERT classifiers trained and tested at the segment level could not perform as well as shown by the performance metrics reported in [23]. Considering that the way how privacy policies are written has not substantially changed over the years, we believe the significantly reduced performances are caused by the inappropriate way of splitting the training and test sets at the segment level. This calls the need to consider the document-level splitting approach for evaluating performances of such classifiers. This is one of the main foci of our work reported later.
IV Extended GDPR Taxonomy and New Corpus

### v2

@@ Abstract
Machine learning (ML) based classifiers that take a privacy policy as the input and predict relevant concepts are useful in different applications such as (semi-)automated compliance analysis against requirements of a specific data protection law such as the EU GDPR. Although many researchers have studied ML-based privacy policy concept classifiers, we observed multiple research gaps, e.g., the lack of a more complete GDPR taxonomy and the less consideration of hierarchical information in privacy policies. To fill such research gaps, we produced a more complete GDPR-oriented privacy policy concept taxonomy, constructed the first privacy policy corpus with explicitly hierarchical information at three levels, and conducted the most comprehensive performance evaluation study of GDPR concept classifiers for privacy policies, cover many aspects that have not been studied systematically. Our work led to multiple findings and insights, including the usefulness of considering hierarchical contextual features and different hierarchical structures, the observation that a “one size fits all” approach may not work, the reduced performance of such classifiers on our newly constructed corpus especially after the first level, and the necessity to split the training and testing sets by documents.
@@ I Introduction
The importance of privacy policies for both data subjects and data controllers means that GDPR-oriented analyses of such legal documents can be very useful. It can help enhance data subjects’ awareness on important data protection issues, and also help data controllers refine their privacy policy to be more legally compliant to the GDPR. Such an analysis can obviously be done qualitatively by experienced GDPR experts, but more automated analysis is preferred in many applications to help reduce the time and costs spent. One frequently studied privacy policy analysis task is concept classification, which involves classifying a given segment of a privacy policy into one or more concepts (see Section II-C for a brief literature review). We observed multiple research gaps in the literature on privacy policy concept classification, including the lack of a more complete GDPR taxonomy and a privacy policy corpus that include more semantic levels and hierarchical information, the lack of consideration of hierarchical information in privacy policies for developing concept classifiers, the lack of studies comparing classifiers with different hierarchical structures, and the inconsistent use of sample splitting methods for generating training and testing sets.
This paper reports our work which helps fill the above research gaps with the following key contributions (ordered following the logical sequence of the contributions).
@@
1) We produced an extended taxonomy for facilitating GDPR-oriented analysis of privacy policies. This taxonomy is based on a smaller GDPR taxonomy proposed by Torre et al. [50], our own knowledge on the GDPR as cyber law experts, two important documents from the ICO and the IAPP (International Association of Privacy Professionals), and the work of the W3C Data Privacy Vocabularies and Controls CG (DPVCG). The end result is the most comprehensive GDPR privacy policy taxonomy reported in the literature so far.
2) We constructed GoPPC-150, the first fully hierarchically encoded GDPR-oriented privacy policy corpus, and a new privacy policy corpus construction framework we used to construct GoPPC-150. GoPPC-150 includes 150 privacy policies collected from Alexa.com top websites. It includes expert-annotated GDPR concept labels in a hierarchical manner following a newly extended GDPR privacy policy taxonomy (see below), and is structured at the document-level. The inclusion of explicitly encoded hierarchical information in GoPPC-150 allows more context-aware development of GDPR concept classifiers and other relevant tools. The framework developed has a high level of automation to minimize human intervention and can be used to further extend GoPPC-150 and to construct other similar corpora.
3) We conducted the most comprehensive performance evaluation study on GDPR concept classifiers (to the best of our knowledge). Our study goes beyond past research by considering the following aspects: i) two hierarchical architectures of such classifiers – local classifier per node (LCN) and local classifier per parent node (LCPN), not just LCPN in past work; ii) multiple sets of traditional features and contextual features reflecting the hierarchical nature of a privacy policy; iii) different machine learning models, and iv) two different sample splitting methods (document- and segment-level). In addition to being the most comprehensive performance evaluation study, we are the first who considered the hierarchical nature of such classifiers and relevant features and investigated the impact of different sample splitting methods on the classification performance.
4) Our comprehensive performance evaluation study led to a range of new findings and insight. These include the usefulness of considering hierarchical contextual features and different hierarchical structures, the observation that a “one size fits all” approach may not work, the reduced performance of such classifiers on our newly constructed corpus especially after the first level, and the necessity to use the document-level sample splitting method for producing the training and testing sets.
@@
Figure 1 visually shows how different parts of our work and key contributions are organized, with the components highlighted in red indicating the new aspects investigated in our work. To help other researchers reproduce our work reported in this paper, all our source code, data used and newly produced, and all results of our experiments have been uploaded to a GitHub repository available at https://github.com/tp-sh/GDPR_privacy_policies. We will refer to this GitHub repository frequently in the rest of the paper, with a direct URL pointing to the corresponding part of the repository when necessary.
Figure 1: A diagram showing different parts of our work reported in this paper.
@@
The rest of this paper is structured as follows. Section II discusses related work. Section III presents the new GDPR-oriented taxonomy and the new privacy policy corpus we developed, detailing the methodology used in their development. Section IV presents our comprehensive performance evaluation of GDPR concept classifiers using our new privacy policy corpus, covering experimental settings and results. Section V provides further discussions of our work and also indicates some future research directions. The last section concludes the paper.
@@ II Related Work
In this section, we introduce related work in four closely related areas: (1) hierarchical multi-label text classification, (2) privacy policy corpora, (3) privacy policy concept classifiers, and (4) analyses of privacy policies. How our work is compared with some selected related work can be found in Table I.
TABLE I: Our work compared with related work (* = new public corpus; = GDPR-specific; texts in bold face = key contributions of our work). The tuples in the “Taxonomy” column indicate the numbers of levels and nodes of the taxonomies, e.g., (3, 96) refers to a taxonomy with 96 nodes at 3 levels. The tuples in the “Corpus/Corpora” column indicate the number of privacy policies, the number of annotation levels and if the corpus includes hierarchical information, e.g., (150, 3, Yes) means that there are 150 privacy policies with annotations at 3 levels, and hierarchical information is included in the annotations. Note that when counting the number of levels, the root node is not counted.
@@
Related Work Taxonomy Features Corpus/Corpora Classifier(s) Sample Splitting Method(s)
OPP-115 [55] (2, 46) Paragraph2Vec OPP-115* (115, 1, No) Multiple; LCPN Document-Level
Polisis [20] No fastText OPP-115 CNN; LCPN Document-Level
Mustapha et al.’s [33] No XLNet OPP-115 XLNet; LCPN Segment-Level
Mousavi Nejad et al.’s [31] No fastText+BERT OPP-115 CNN+BERT; LCPN Segment-Level
PrivBERT [47] No RoBERTa OPP-115 CNN; LCPN Segment-Level
Torre et al.’s [3] (3, 69) GloVe Private SVM; LCN Document-Level
Rahat et al.’s [2] (1, 18) fastText Private CNN; LCPN Document-Level
Our Work (3, 96) Multiple; Context GoPPC-150* (150, 3, Yes) Multiple; LCPN & LCN Document-Level & Segment-Level
II-A Hierarchical Multi-Label Text Classification (HMTC)
Privacy policy is a hierarchical text with one or more labels for each sentence / paragraph, so the concept classification task of a given privacy policy can be regarded as an HMTC task. According to Silla et al. [45], there are three main approaches to HMTC – flat (learning about leaf nodes only so completely ignoring the hierarchical structure), local (learning about each node or sets of nodes separately) and global (learning all nodes at the same time) classifiers. Flat classifiers are naive and global ones are usually too complicated and require sufficient training data for all nodes. Therefore, local classifiers have received the most attention from researchers, and multiple types have been developed, including LCN (a local classifier per node), LCPN (a local classifier per parent node) as well as LCL (a local classifier per level) classifiers [13, 16, 44].
When it comes to privacy policy concept classification, most researchers also paid more attention to local classifiers [20, 33, 50, 2, 47], particularly LCPN classifiers based on CNN. The “Classifier(s)” column of Table I shows the type of classifiers studied in some selected related work. As can be seen, LCN classifiers have been much less studied and we were unaware of any studies comparing performances of LCPN and LCN classifiers. LCL classifiers are usually not appropriate for privacy policy concept classification since concepts belonging to different branches are often too different to be processed by a single classifier. This is why we have not seen any related work using LCL classifiers.
II-B Privacy Policy Corpora
@@
In 2016, Wilson et al. [55] created a public privacy policy corpus called “Online Privacy Policy” (OPP-115) by hiring three legal experts as annotators and it has been used by many other researchers. However, Sarne et al. [42] used unsupervised ML to model topics in privacy policies and found a mismatch between the topics in privacy policies they analyzed and topics covered in OPP-115. Sathyendra et al. [43] constructed a more fine-grained corpus based on OPP-115 with semi-automated annotation, focusing on “opt-out” in privacy policies. Leblanc and Liu [26] built a corpus of fuzzy words and sentences in privacy policies. Zimmeck et al. [64] created the App-350 Privacy Policy Corpus with the goal of checking compliance of behaviors and privacy policies of mobile apps. They selected 350 privacy policies of the most popular apps on Google Play and hired legal experts as annotators. Robaldo et al. [39] used I/O logic formula to code and model various legal documents, expressed legal statements in a logical language, explained various clauses, and demonstrated their work with the GDPR. Muller et al. [32] introduced a privacy policy dataset containing over 18,300 sentences, tagged according to five core privacy policy requirements of the GDPR. Srinath et al. [47] created PrivaSeer, an automatically constructed corpus containing privacy policies of over 1 million English websites, and studied the composition of the corpus, showing readability tests, document similarity, keyword extraction results, and exploring the corpus through topic modeling. Kuznetsov et al. [25] used a technique for identifying URLs of privacy policies of IoT devices to construct a new corpus with 592 privacy policies in 2022. Zhan et al. [60] constructed VRPP, the first large-scale dataset of privacy policies for virtual reality applications, comprising 11,923 distinct VR apps from 10 mainstream platforms.
II-C Privacy Policy Concept Classifiers
@@
Harkous et al. [20] proposed Polisis, a comprehensive framework for enabling multi-dimensional privacy policy analysis. Polisis provides an online service for privacy policies analysis, using a combination of NLP and deep learning (DL) techniques to extract fragments from privacy policies, each containing a set of labels describing data processing behaviors. Tesfay et al. [48] proposed an ML-based approach to classify multiple categories of privacy policy content using pre-defined keywords. Lippi et al. [28] provided 33 metadata types for privacy policies in terms of the GDPR compliance, and provided automatic support for ambiguity detection of privacy policies based on manual rules and machine learning (ML) based methods. A more comprehensive piece of work was done by Torre et al. [50], who proposed 55 metadata types, covering all types identified by Lippi et al. [28], and studied automated detection of privacy policy integrity using an advanced combination of natural language processing (NLP) and ML based on the 55 metadata types. Their more recent work [3] further improved their privacy policy completeness checking framework with the same GDPR taxonomy and classifiers. Mousavi Nejad et al. [31] studied the performance of multiple machine learning techniques (CNN and two BERT-based transformers) on classifying privacy privacy concepts in the OPP-115 corpus, and reported the best performance figures of 77–85% (micro-avg) and 76–79% (macro-avg). Mustapha et al. [33] proposed to use a pre-trained XLNet and a dense layer for classification of concepts in OPP-115, and showed the classifier can outperform the best models reported in [31] by 1-3% for both macro and micro average F1-scores. Srinath et al. [47] constructed a new privacy policy corpus called PrivSeer, built a pre-trained language model called PrivBERT based on PrivSeer, and then developed a number of concept classifiers for privacy policy analysis based on PrivBERT. As a whole, the best-performing classifiers are PrivBERT-based ones and those reported in [3], achieving an F1 score above 0.8 for most concept classifiers, and even above 0.9 for some. Xie et al. [58] constructed a collection of privacy policy clauses grounded in multiple privacy regulations, leveraging self-annotated excerpts from both open-source datasets and real-world policies, and applied zero-shot prompting with large language models for clause classification.
II-D Analyses of Privacy Policies
In addition to privacy policy concept classification, there are many other past studies focusing on automated or semi-automated analyses of privacy policies, some of which are GDPR-oriented. Many of such studies are based on machine learning based concept classifiers, as part of a larger automated system.
@@ II-D Analyses of Privacy Policies
When it comes to the GDPR-oriented analyses, Tom et al. [49] presented a preliminary GDPR model that aims to provide a simple and visual overview for human operators to achieve a better understanding of the relationships between different concepts in the GDPR. It also describes a method of using their proposed model as a tool to develop privacy policies and illustrates how to extract compliance rules. Palmirani and Governatori [35] proposed a proof of concept applicable to the GDPR domain with the aim of detecting or preventing violations of privacy enforcement norms. Bhatia et al. [5] identified incompleteness of privacy policies by representing data practice descriptions as semantic frames. The approach was a grounded analysis to discover which semantic roles corresponding to a data action are needed to construct complete data practice descriptions. Mousavi et al. [34] proposed a tool called KnIGHT, whose innovation lies in the use of semantic similarity between words to associate sentences in a privacy policy with relevant paragraphs in the legal text of the GDPR. Torre et al. [51] proposed a model-based solution for GDPR-compliance analysis, using Unified Modeling Language (UML) and Object Constraint Language (OCL) to build a UML representation of the GDPR. Hamdani et al. [19] conceptualized a framework to implement document-central compliance checking methods in the data supply chain and developed concrete methods to automatically check GDPR-compliance of privacy policies. Xiang et al. [57] developed PolicyChecker, a rule-based framework that encodes the mandatory and context-dependent requirements from GDPR Articles 13 and 14 into logical chains to automatically assess the completeness of mobile app privacy policies, evaluated on 205,973 Android apps from the UK Google Play Store.
Some other work focused on the commonalities of privacy policies. In order to increase transparency of privacy policies, Zimmeck et al. [62] proposed Privee, a software architecture for analyzing essential policy terms based on crowd sourcing and automatic classification techniques. They implemented Privee as a proof-of-concept web browser extension that retrieves policy analysis results from an online privacy policy repository or, if no such results are available, performs automatic classifications. Targeting privacy policies of websites and mobile apps, Caramujo et al. [10] proposed a domain-specific language and a model transformation approach for specifying privacy policy models. Pullonen et al. [38] proposed a multi-level model as an extension of the business process model and representation to support visualization, analysis, and communication of the privacy policy characteristics of business processes. Ayala-Rivera and Pasquale [4] proposed a model-based approach to help online services understand their data protection obligations under the GDPR. Zimmeck et al. [63] designed and implemented PrivacyFlash Pro, an automated privacy policy generator for iOS apps that leverages static analysis, which identifies code signatures composed of Plist permission strings, framework imports, class instantiations, authorization methods and other evidence that are mapped to privacy practices expressed in privacy policies. Zhou et al. [61] proposed POLICYCOMP, an automated framework that identifies overbroad personal data collection practices in mobile app privacy policies by comparing them against functionally similar counterpart apps, revealing that nearly half of such practices may exceed actual service needs. Wang et al. [53] proposed PrivGuard, a novel system design that reduces human participation required and improves the productivity of the compliance process, which is mainly comprised of two components: (1) PrivAnalyzer, a static analyzer based on abstract interpretation for partly enforcing privacy regulations, and (2) a set of components providing strong security protection on the data throughout its life cycle.
Some work has a focus on helping users to better comprehend privacy policies, especially via visual and graph-based mechanisms. Cui et al. [14] proposed PoliGraph, automated analysis of the information disclosed in a policy into a knowledge graph based on NLP technology. Although they also combine the context of the current text for classification, they only focus on the data collection part of the privacy policy, and do not consider the structural information of the privacy policy. To enhance user understanding of data handling implications, Zimmermann et al. [65] conducted a between-subjects study, evaluating the effectiveness of visual metaphors and dynamic feedback in supporting informed privacy decisions. Pan et al. [36] proposed SeePrivacy, a multimodal framework that combines GUI visual analysis and large language models to extract context-relevant segments from privacy policies in mobile applications, aiming to enhance user interaction with and understanding of those policies. Adhikari et al. [1] introduced PolicyPulse, an NLP pipeline that focuses on extracting semantic roles of different data entities from a privacy policy to enhance users’ comprehension of privacy policies. The role analysis is achieved by first classifying different semantic frames into relevant data practice categories, and then relevant frames are further processed to extract semantic roles. The frame categorization step is very similar to the concept classification discussed in the previous subsection, but PolicyPulse only involves 5 selected categories from OPP-115 that are more useful for supporting the semantic role extraction. This work therefore is not directly about concept classification in privacy policies. Yang et al. [59] proposed a framework that uses large language models and knowledge distillation to automatically extract structured privacy attributes from policies and represent them as queryable graphs.
III Extended GDPR Taxonomy and New Corpus